Show that for one mole of a perfect gas, π‘‘π‘†π‘š = 𝐢𝑃,π‘š

𝑑𝑇

𝑇

βˆ’ 𝑅

𝑃

𝑑


1
Expert's answer
2021-03-16T08:25:07-0400

Entropy is by its definition: dS = (dU+PdV)T

Multiply that by T: TdS - PdV = dU = d(H-PV) = dH - PdV - VdP

So: TdS = dH - VdP

For an ideal gas: PV=nRT or PVm=RT, dU = cvdT, dH = cpdT

Finally, we get dSm = cp,mdT/T - RdP/P